Find the intersection of the line and plane:

3y−x+z=−4,r(t)=⟨−1,2,1⟩+t⟨−2,−3,1⟩

3y−x+z=−4,r(t)=⟨−1,2,1⟩+t⟨−2,−3,1⟩

P=(

P=(  , , )


1
Expert's answer
2022-01-19T16:26:21-0500

"r(t)"

"=\u27e8-1,2,1\u27e9+t\u27e8-2,-3,1\u27e9"

"=\u27e8-1,2,1\u27e9+\u27e8-2t,-3t,t\u27e9"

Adding, this gives

"=\u27e8-1-2t,2-3t,1+t\u27e9"

"\implies x=-1-2t"

"y=2-3t"

"z=1+t"


Substituting "x,y,z" into "3y-x+z", we have

"3(2-3t)-(-1-2t)+(1+t)=-4"

"\implies6-9t+1+2t+1+t=-4"

"\implies-9t+2t+t+6+1+1=-4"

"\implies-6t+8=-4"

"\implies-6t=-4-8"

"\implies-6t=-12"

Dividing both sides by "-6" .This gives

"t=\frac{-12}{-6}=2"


Substituting "t" into "\u27e8-1-2t,2-3t,1+t\u27e9"

We have

"\u27e8\u22121\u22122(2),2\u22123(2),1+2\u27e9"

"P=\u27e83,-4,3\u27e9"